OpenStudy (cutiecomittee123):

I am so confused with this question, someone please lead me to an understanding "There are values of t so that cos t= 5/3"

OpenStudy (compassionate):

Huh? There are values of t, so that cos(t) = 5/3 I think what it's saying is, cos(t) [Some number] is equal to 5/3 In other words, what number of cos will be 5/3, or 1.6666....

OpenStudy (cutiecomittee123):

well the choices given are true or false so it is a true or false question

OpenStudy (anonymous):

lol the answer is NO!!

OpenStudy (anonymous):

since cosine cannot be larger than 1

OpenStudy (compassionate):

\[\cos(t) = 1.67 \space \space \space or \space \space \space \frac{ 5 }{ 3 }\] I guess you could plug in terms, like cos(1), cos(2), and if you can't find the value, then it's false.

OpenStudy (cutiecomittee123):

okay cool that makes sense then

OpenStudy (anonymous):

Not possible. There is no value greater than 1 on the unit circle

OpenStudy (anonymous):

the largest cosine can be is 1, the smallest is -1

OpenStudy (compassionate):

@satellite73 , let her learn, my child. Let her learn.

OpenStudy (zzr0ck3r):

let her learn by plugging in infinite values until she reaches the conclusion?

OpenStudy (zzr0ck3r):

By definition, cos(x), sin(x) live between -1 and 1

OpenStudy (cutiecomittee123):

okay so no because cos cannot be bigger than 1

OpenStudy (zzr0ck3r):

correct